OpenStudy (anonymous):

find the measure of an interior angle of a regular polygon with 9 sides

OpenStudy (anonymous):

SUM of all interior angles = 180(n-2) with n = 9 sides, SUM = 180(9-2) = 1260 So each interior angle of this regular 9-sided polygon will be 1260/9 = 140 degrees.

OpenStudy (anonymous):

Make sure you understand what is meant by a REGULAR polygon.

OpenStudy (anonymous):

thank you soooo much!!!! do you think you can help me with some others that i have?

OpenStudy (anonymous):

Welcome.

OpenStudy (anonymous):

are you good with point slope things?